# Ensemble Kalman Diffusion Guidance: A Derivative-free Method for Inverse Problems

### v1

@@ 3.2 Our approach: Ensemble Kalman Diffusion Guidance
6:     $g^{(j)}_{i}\leftarrow\frac{1}{J}\sum_{k=1}^{J}\left\langle G(\hat{{\bm{x}}}^{(k)}_{N})-\bar{G},{\bm{y}}-G(\hat{{\bm{x}}}^{(j)}_{N})\right\rangle_{\Gamma}\left({\bm{x}}^{(j)}_{i}-\bar{{\bm{x}}_{i}}\right)$
@@
where $\hat{{\bm{x}}}_{N}$ is obtained by running the Probability Flow ODE solver $\phi$ starting at ${\bm{x}}_{i}$. One attractive property of this estimate compared to popular ones based on $\mathbb{E}[{\bm{x}}_{N}|{\bm{x}}_{i}]$ and isotropic Gaussian approximations in previous works Chung et al. (2022a); Song et al. (2023a; b) is that our approximation stays on data manifold but the Gaussian approximations include additive noise that do not live on data manifold. This aspect is particularly important for scientific inverse problems based on partial differential equations (PDEs), where staying on the data manifold is important for reliably solving the forward model $p({\bm{y}}|{\bm{x}})$. For instance, we observe that Gaussian approximations frequently violate the stability conditions of numerical PDE solvers, leading to meaningless estimates.
@@
Consider an ensemble of particles $\{{\bm{x}}_{i}^{(j)}\}_{j=1}^{J}$. Let $\bar{{\bm{x}}}_{i}$ denote their empirical mean and ${\bm{C}}_{xx}^{(i)}$ denote their empirical covariance matrix, at the $i$-th iteration.
$\bar{{\bm{x}}}_{i}=\frac{1}{J}\sum_{j=1}^{J}{\bm{x}}_{i}^{(j)},{\bm{C}}_{xx}^{(i)}=\frac{1}{J}\sum_{j=1}^{J}\left({\bm{x}}_{i}^{(j)}-\bar{{\bm{x}}}_{i}\right)\left({\bm{x}}_{i}^{(j)}-\bar{{\bm{x}}}_{i}\right)^{\top}.$
Instead of the commonly used scalar weight $w_{i}$, we use a weighting matrix $w_{i}{\bm{C}}_{xx}^{(i)}$ in Eq. (8):
${\bm{x}}_{i+1}^{(j)}\approx\arg\min_{{\bm{x}}_{i+1}^{(j)}}$ $\frac{1}{2}\left({\bm{x}}_{i+1}-{\bm{x}}^{\prime}_{i}\right)^{\top}\left(w_{i}{\bm{C}}_{xx}^{(i)}\right)^{-1}\left({\bm{x}}_{i+1}-{\bm{x}}^{\prime}_{i}\right)$ (12)
$-\nabla^{\top}\log\hat{p}\left({\bm{y}}|{\bm{x}}_{i}^{\prime(j)}\right)\left({\bm{x}}_{i+1}-{\bm{x}}_{i}^{\prime(j)}\right)$ (13)
$={\bm{x}}_{i}^{\prime(j)}+$ $w_{i}{\bm{C}}_{xx}^{(i)}\nabla\log\hat{p}\left({\bm{y}}|{\bm{x}}_{i}^{\prime(j)}\right),$ (14)
which can be viewed as a gradient step preconditioned by ${\bm{C}}_{xx}^{(i)}$. Leveraging the idea of statistical linearization in the ensemble Kalman methods (Bergemann & Reich, 2010; Schillings & Stuart, 2017), we can perform the preconditioned gradient step without explicit derivative.
@@
Under Assumption 12 and 3, suppose the algorithm is implemented with $w_{i}=1/\left(tr\left(C_{yy}^{(i)}\right)\right)$, then we have the following approximation:
${\bm{C}}_{xx}^{(i)}\nabla\log\hat{p}\left({\bm{y}}|{\bm{x}}_{i}^{\prime(j)}\right)\approx\frac{1}{J}\sum_{k=1}^{J}\left\langle\psi\left({\bm{x}}^{\prime(k)}_{i}\right)-\bar{G},{\bm{y}}-\psi\left({\bm{x}}^{\prime(j)}_{i}\right)\right\rangle_{\Gamma}\left({\bm{x}}^{\prime(j)}_{i}-\bar{{\bm{x}}_{i}}\right),$ (16)
@@
The detailed derivation can be found in Appendix A.2. Therefore, we can update each particle with the ensemble difference to move toward the high-likelihood region without evaluating the derivative. A full description of our method is given in Algorithm 2. We name it Ensemble Kalman Diffusion Guidance as it is a guidance for diffusion models based on ensemble Kalman approximation. Implementation details are provided in Appendix A.4.
@@ 4 Experiments
We empirically study our method on the classic image restoration problems and two scientific inverse problems. We view the scientific inverse problems as the more interesting domains for evaluating our method, particularly the Navier-Stokes equation where it is impractical to accurately compute the gradient of the forward model.
@@
We focus on comparing against methods that only use black-box access to the forward model. Our first two baselines incorporate numerical estimation methods to substitute automatic differentiation, named Forward-GSG and Central-GSG given in Algorithm 3. Specifically, we use forward Gaussian smoothed gradient (Forward GSG in Eq. 28) and central Gaussian smoothed gradient (Central GSG in Eq. 29) to approximate the gradient of the log-likelihood, and plug it into the standard gradient-based method, Diffusion Posterior Sampling (DPS)(Chung et al., 2023b). more details can be found in Appendix A.3. The third baseline is Stochastic Control Guidance (SCG) proposed in Huang et al. (2024), which is designed to deal with non-differentiable guidance. The last baseline in Diffusion Policy Gradient (DPG) proposed by Tang et al. (2024). Additionally, for Navier-Stokes equation, we add the traditional method Ensemble Kalman Inversion (EKI) proposed in Iglesias et al. (2013).
Table 1: Quantitative evaluation on FFHQ 256x256 dataset. We report average metrics for image quality and samples consistency on four tasks. Measurement noise level $\sigma=0.05$ is used if not otherwise stated.
@@
Table 2: Comparison of different derivative-free algorithms on Navier-Stokes inverse problem. Numbers in parentheses represent the sample standard deviation.
Relative L2 Relative L2 Relative L2 Total # FME Total # DME Seq # FME Seq # DME
EKI 0.577 (0.138) 0.609 (0.119) 0.673 (0.107) 1024k 0 0.50k 0
Forward-GSG 1.687 (0.156) 1.612 (0.173) 1.454 (0.154) 2049k 1k 1k 1k
Central-GSG 2.203* (0.314) 2.117 (0.295) 1.746 (0.191) 2048k 1k 1k 1k
DPG 0.325 (0.188) 0.408* (0.173) 0.466 (0.171) 4000k 1k 1k 1k
SCG 0.908 (0.600) 0.928 (0.557) 0.966 (0.546) 512k 384k 0.75k 1k
EnKG(Ours) 0.120 (0.085) 0.191 (0.057) 0.294 (0.061) 295k 2695k 0.14k 1.3k
4.1 Image inverse problems
Tackling image inverse problems (e.g., deblurring) is common in the literature and serves as a reasonable reference domain for evaluation. We note that we consider a harder version of the problem where we do not use the gradient of the forward model. Moreover, most imaging problems use a linear forward model (except for phase retrieval), which is significantly simpler than the nonlinear forward models more often used in scientific domains. We do note that, perhaps surprisingly, for the nonlinear phase retrieval task, our approach outperforms the standard strong DPS baseline which actually uses the gradient of the forward model.
Problem setting
We evaluate our algorithm on the standard image inpainting, superresolution, deblurring (Gaussian), and phase retrieval problems. For inpainting, the forward model is a box mask with randomized location. For superresolution, we employ bicubic downsampling (either ×$\times 2$ or ×$\times 4$) and for Gaussian deblurring, a blurring kernel of size $61\times 61$ with standard deviation 3.0. Finally, phase retrieval takes the magnitude of the Fourier transform of the image as the observation. We use measurement noise $\sigma=0.05$ in all experiments except for superresolution on $64\times 64$ images, where we set $\sigma=0.01$. The pre-trained diffusion model for FFHQ $64\times 64$ is taken unmodified from Karras et al. (2022). The model for FFHQ $256\times 256$ is taken from Chung et al. (2022a) and converted to the EDM framework using their Variance-Preserving (VP) preconditioning.
@@ Results
We demonstrate the qualitative results for image inverse problems in Figure 7, and show the quantitative results in Table 5 and 1. EnKG achieves comparable or even better performance compared to DPS (Chung et al., 2022a), which has gradient access to the forward function and the diffusion model. EnKG consistently outperforms all baseline methods with black-box access on two data settings and four different tasks.
4.2 Navier Stokes
@@
The goal is to recover the initial vorticity field from a noisy sparse observation of the vorticity field at time $T=1$. The Navier-Stokes equation  17 does not admit a closed form solution and thus there is no closed form derivative available for the solution operator. Moreover, obtaining an accurate numerical derivative via automatic differentiation through the numerical solver is challenging due to the extensive computation graph expanded after thousands of discrete time steps.
We first solve the equation up to time $T=5$ using initial conditions from a Gaussian random field, which is highly complicated due to the nonlinearity of Navier-Stokes equation. We sample 20,000 vorticity fields from this empirical measure to train our diffusion model. Then, we independently sample 10 random vorticity fields as the test set.
@@
We report the relative $L^{2}$ error to evaluate the accuracy of the algorithm, which is given by $\frac{\left\|\hat{{\bm{w}}}_{0}-{\bm{w}}^{*}_{0}\right\|_{L^{2}}}{\left\|{\bm{w}}^{*}_{0}\right\|_{L^{2}}}$ where $\hat{{\bm{w}}}_{0}$ is the predicted vorticity and ${\bm{w}}^{*}_{0}$ is the ground truth. To comprehensively capture the computational characteristics of different algorithms, we use the following metrics: 1) Total # Fwd: total number of forward model evaluations; 2) Seq. # Fwd: number of sequential forward model evaluations; 3) Total # DM: total number of diffusion model evaluations; 4) Seq. # DM: number of sequential diffusion model evaluations; 5) Total # DM grad: total number of diffusion model gradient evaluations; 6) Seq. # DM grad: number of sequential diffusion model gradient evaluations.
@@ Results
In Table 2, we show the average relative $L^{2}$ error of the recovered ground truth at different noise levels of the observations. Our approach, by far, outperforms methods which use other forms of approximate derivative information. Indeed, we show in Figure 2 that our method give solutions which qualitatively preserve important features of the flow while other methods completely fail to give a reasonable answer. Additionally, we compare the results in the radial power spectrum as shown in Figure 3(a). The traditional baseline EKI produces an over-smoothed solution, while DPG,SCG, and GSG tend to generate noisy solutions. In Figure 3 (b), we visualize the main source of computation costs, showing that the forward model evaluation of Navier-Stokes equation is significantly slower than other parts as the numerical solver has to take thousands of iterations. Figure 3 (c) compares different algorithms with our proposed efficiency metrics, showing that EnKG requires the least number of forward model evaluations and is highly parallelizable. Therefore, for problems where the forward model evaluation is slow, EnKG stands out to be the most efficient algorithm compared to the baseline methods.
Figure 3: Panel (a): radial power spectrum of the solutions from different algorithms. Panel (b): runtime of single evaluation of the forward model, diffusion model, and diffusion model gradient (tested on a single A100). Panel (c): comparison of computational characteristics of different algorithms on Navier-Stokes problem.

### v2

@@ 4.2 Our approach: Ensemble Kalman Diffusion Guidance
6:    $g^{(j)}_{i}\leftarrow\frac{1}{J}\sum_{k=1}^{J}\left\langle G(\hat{{\bm{x}}}^{(k)}_{N})-\bar{G},{\bm{y}}-G(\hat{{\bm{x}}}^{(j)}_{N})\right\rangle_{\Gamma}\left({\bm{x}}^{(k)}_{i}-\bar{{\bm{x}}_{i}}\right)$
@@
where $\hat{{\bm{x}}}_{N}$ is obtained by running the PF ODE solver $\phi$ starting at ${\bm{x}}_{i}$. One attractive property of this estimate compared to popular ones based on $\mathbb{E}[{\bm{x}}_{N}|{\bm{x}}_{i}]$ and isotropic Gaussian approximations in previous works Chung et al. (2022a); Song et al. (2023a; b) is that our approximation stays on data manifold but the Gaussian approximations include additive noise that do not live on data manifold. This aspect is particularly important for scientific inverse problems based on partial differential equations (PDEs), where staying on the data manifold is important for reliably solving the forward model $p({\bm{y}}|{\bm{x}})$. For instance, we observe that Gaussian approximations frequently violate the stability conditions of numerical PDE solvers, leading to meaningless estimates.
@@
Consider an ensemble of particles $\{{\bm{x}}_{i}^{(j)}\}_{j=1}^{J}$. Let $\bar{{\bm{x}}}_{i}$ denote their empirical mean and $C_{xx}^{(i)}$ denote their empirical covariance matrix, at the $i$-th iteration:
$\bar{{\bm{x}}}_{i}=\frac{1}{J}\sum_{j=1}^{J}{\bm{x}}_{i}^{(j)},\quad C_{xx}^{(i)}=\frac{1}{J}\sum_{j=1}^{J}\left({\bm{x}}_{i}^{(j)}-\bar{{\bm{x}}}_{i}\right)\left({\bm{x}}_{i}^{(j)}-\bar{{\bm{x}}}_{i}\right)^{\top}.$
Instead of the commonly used scalar weight $w_{i}$, we use a weighting matrix $w_{i}C_{xx}^{(i)}$ in Eq. (8):
${\bm{x}}_{i+1}^{(j)}\approx\arg\min_{{\bm{x}}_{i+1}}$ $\frac{1}{2}\left({\bm{x}}_{i+1}-{\bm{x}}^{\prime(j)}_{i}\right)^{\top}\left(w_{i}C_{xx}^{(i)}\right)^{-1}\left({\bm{x}}_{i+1}-{\bm{x}}^{\prime(j)}_{i}\right)$ (12)
$-\nabla_{{\bm{x}}}^{\top}\log\hat{p}\left({\bm{y}}|{\bm{x}}_{i}^{\prime(j)}\right)\left({\bm{x}}_{i+1}-{\bm{x}}_{i}^{\prime(j)}\right)$ (13)
$={\bm{x}}_{i}^{\prime(j)}+$ $w_{i}C_{xx}^{(i)}\nabla_{{\bm{x}}}\log\hat{p}\left({\bm{y}}|{\bm{x}}_{i}^{\prime(j)}\right).$ (14)
Note that in practice, $C_{xx}^{(i)}$ can be singular when the number of particles is smaller than the particle dimension. In such cases, the matrix inverse in Eq. (12) is generalized to the sense of the Moore-Penrose inverse as $C_{xx}^{(i)}$ is always positive semi-definite. Eq. (14) effectively becomes a gradient projected onto the subspace spanned by the particles. At its current form, Eq. (14) still requires the gradient information. Next, we show how to approximate this gradient step without explicit derivative by Leveraging the idea of statistical linearization in the ensemble Kalman methods (Bergemann & Reich, 2010; Schillings & Stuart, 2017).
@@
Remark
To verify that Assumption 3 holds in most cases of interest, we plot the traces $tr\left(C_{xx}\right)$ and $tr\left(C_{yy}\right)$ across three experiments of different inverse problems. As shown in Figure 5, while larger $tr\left(C_{xx}\right)$ does not always indicate larger $tr\left(C_{yy}\right)$ due to the ill-posedness, $tr(C_{yy})$ approaches zero only when $tr(C_{xx})$ also approaches zero, aligning with our assumption.
@@
Under Assumption 12 and 3, suppose the correction step is implemented as follows with $w_{i}=1/\left(tr\left(C_{yy}^{(i)}\right)\right)$,
${\bm{x}}_{i+1}^{(j)}$ $={\bm{x}}_{i}^{\prime(j)}+w_{i}C_{xy}^{(i)}\left({\bm{y}}-\psi\left({\bm{x}}_{i}^{\prime(j)}\right)\right)$ (16)
$={\bm{x}}_{i}^{\prime(j)}+w_{i}\frac{1}{J}\sum_{k=1}^{J}\left\langle\psi\left({\bm{x}}^{\prime(k)}_{i}\right)-\bar{G},{\bm{y}}-\psi\left({\bm{x}}^{\prime(j)}_{i}\right)\right\rangle_{\Gamma}\left({\bm{x}}^{\prime(k)}_{i}-\bar{{\bm{x}}_{i}}\right),$ (17)
where
$C_{xy}^{(i)}=\frac{1}{J}\sum_{j=1}^{J}\left({\bm{x}}_{i}^{\prime(j)}-\bar{{\bm{x}}}_{i}\right)\left(\psi\left({\bm{x}}_{i}^{\prime(j)}\right)-\bar{\psi}_{i}\right)^{\top}.$
After sufficient iterations, we have the following approximation:
$C_{xy}^{(i)}\left({\bm{y}}-\psi\left({\bm{x}}_{i}^{\prime(j)}\right)\right)$ $=\frac{1}{J}\sum_{k=1}^{J}\left\langle\psi\left({\bm{x}}^{\prime(k)}_{i}\right)-\bar{G},{\bm{y}}-\psi\left({\bm{x}}^{\prime(j)}_{i}\right)\right\rangle_{\Gamma}\left({\bm{x}}^{\prime(k)}_{i}-\bar{{\bm{x}}_{i}}\right)$ (18)
$\approx C_{xx}^{(i)}\nabla_{{\bm{x}}}\log\hat{p}\left({\bm{y}}|{\bm{x}}_{i}^{\prime(j)}\right),$ (19)
@@
The detailed derivation can be found in Appendix A.2. Proposition 1 shows that the ensemble update step defined in Eq. (17) effectively approximates the preconditioned gradient step defined in Eq. (12) without explicit derivative:
${\bm{x}}_{i+1}^{(j)}={\bm{x}}_{i}^{\prime(j)}+w_{i}C_{xy}^{(i)}\left({\bm{y}}-\psi\left({\bm{x}}_{i}^{\prime(j)}\right)\right)\approx{\bm{x}}_{i}^{\prime(j)}+w_{i}C_{xx}^{(i)}\nabla_{{\bm{x}}}\log\hat{p}\left({\bm{y}}|{\bm{x}}_{i}^{\prime(j)}\right).$ (20)
Algorithm 2 puts it all together and provides a complete description of the proposed method. Implementation details are provided in Appendix A.4.
@@ 5 Experiments
We empirically study our EnKG method on the classic image restoration problems and two scientific inverse problems. We view the scientific inverse problems as the more interesting domains for evaluating our method, particularly the Navier-Stokes equation where it is impractical to accurately compute the gradient of the forward model.
@@
We focus on comparing against methods that only use black-box access to the forward model. The first two baselines, Forward-GSG and Central-GSG (Algorithm 3), use numerical estimation methods instead of automatic differentiation to approximate the gradient of the log-likelihood, and plug it into a standard gradient-based method, Diffusion Posterior Sampling (DPS) (Chung et al., 2023b). Specifically, Forward-CSG uses a forward Gaussian smoothed gradient (Eq. 37), and Central-CSG uses a central Gaussian smoothed gradient (Eq. 38). More details are in Appendix A.3. The last two baselines are Stochastic Control Guidance (SCG) (Huang et al., 2024) and Diffusion Policy Gradient (DPG) (Tang et al., 2024), discussed in Sec. 2. For Navier-Stokes, we also add the conventional Ensemble Kalman Inversion (EKI) (Iglesias et al., 2013).
5.1 Image inverse problems
Tackling image inverse problems (e.g., deblurring) is common in the literature and serves as a reasonable reference domain for evaluation. We note that we consider a harder version of the problem where we do not use the gradient of the forward model. Moreover, most imaging problems use a linear forward model (except for phase retrieval), which is significantly simpler than the non-linear forward models more often used in scientific domains.
Problem setting
We evaluate our algorithm on the standard image inpainting, superresolution, deblurring (Gaussian), and phase retrieval problems. For inpainting, the forward model is a box mask with randomized location. For superresolution, we employ bicubic downsampling (either ×$\times 2$ or ×$\times 4$) and for Gaussian deblurring, a blurring kernel of size $61\times 61$ with standard deviation 3.0. Finally, phase retrieval takes the magnitude of the Fourier transform of the image as the observation. We use measurement noise $\sigma=0.05$ in all experiments except for superresolution on $64\times 64$ images, where we set $\sigma=0.01$. The pre-trained diffusion model for FFHQ $64\times 64$ is taken unmodified from Karras et al. (2022). The model for FFHQ $256\times 256$ is taken from Chung et al. (2022a) and converted to the EDM framework (Karras et al., 2022) using their Variance-Preserving (VP) preconditioning.
Table 1: Quantitative evaluation on FFHQ 256x256 dataset. We report average metrics for image quality and consistency on four tasks. Measurement noise is $\sigma=0.05$ unless otherwise stated.
@@ Results
We show the quantitative results in Table 1(Appendix), and qualitative results in Figure 7 (Appendix). On the easier linear inverse problems (inpainting, superresolution, and deblur), EnKG comes in second to DPG. On the harder non-linear phase retrieval problem, EnKG is comfortably the best approach. This trend is consistent with our results in the scientific inverse problems, which are all non-linear.
5.2 Navier-Stokes equation
@@
The goal is to recover the initial vorticity field from a noisy sparse observation of the vorticity field at time $T=1$. Eq. (21) does not admit a closed form solution and thus there is no closed form derivative available for the solution operator. Moreover, obtaining an accurate numerical derivative via automatic differentiation through the numerical solver is challenging due to the extensive computation graph that can span thousands of discrete time steps. We first solve the equation up to time $T=5$ using initial conditions from a Gaussian random field, which is highly complicated due to the non-linearity of Navier-Stokes equation. We sample 20,000 vorticity fields to train our diffusion model. Then, we independently sample 10 random vorticity fields as the test set.
Table 2: Comparison on the Navier-Stokes inverse problem. Numbers in parentheses represent the sample standard deviation. Metrics to evaluate computation costs are defined in Sec. 5.2. *: one or two test cases are excluded from the results due to numerical instability. Runtime is reported on a single A100 GPU.
Relative L2 Relative L2 Relative L2 Total # Fwd Total # DM Seq # Fwd Seq # DM Runtime
EKI 0.577 (0.138) 0.609 (0.119) 0.673 (0.107) 1024k 0 0.50k 0 121 mins
Forward-GSG 1.687 (0.156) 1.612 (0.173) 1.454 (0.154) 2049k 1k 1k 1k 105 mins
Central-GSG 2.203* (0.314) 2.117 (0.295) 1.746 (0.191) 2048k 1k 1k 1k 105 mins
DPG 0.325 (0.188) 0.408* (0.173) 0.466 (0.171) 4000k 1k 1k 1k 228 mins
SCG 0.908 (0.600) 0.928 (0.557) 0.966 (0.546) 384k 384k 0.75k 1k 119 mins
EnKG(Ours) 0.120 (0.085) 0.191 (0.057) 0.294 (0.061) 295k 3342k 0.14k 1.3k 124 mins
@@
We report the relative $L^{2}$ error to evaluate the accuracy of the algorithm, which is given by $\frac{\left\|\hat{{\bm{w}}}_{0}-{\bm{w}}^{*}_{0}\right\|_{L^{2}}}{\left\|{\bm{w}}^{*}_{0}\right\|_{L^{2}}}$ where $\hat{{\bm{w}}}_{0}$ is the predicted vorticity and ${\bm{w}}^{*}_{0}$ is the ground truth. To comprehensively analyze the computational requirements of inverse problem solvers, we use the following metrics: the total number of forward model evaluations (Total # Fwd); the number of sequential forward model evaluations (Seq. # Fwd), where each evaluation depends on the previous one.; the total number of diffusion model evaluations (Total # DM); the number of sequential diffusion model evaluations (Seq. # DM), which is analogous to Seq. # Fwd but focuses on diffusion model evaluation; the total number of diffusion model gradient evaluations (Total # DM grad); the number of sequential diffusion model gradient evaluations (Seq. # DM grad). These metrics are designed to reflect the primary computational demands: forward model queries and diffusion model queries. Sequential metrics are particularly important because they determine the minimum runtime achievable, independent of the available computational resources. By isolating sequential evaluations, we can better assess the parallelization potential of the algorithm, akin to the “critical path” concept in algorithm analysis from the computer science literature (Kohler, 1975).
@@ Results
In Table 2, we show the average relative $L^{2}$ error of the recovered ground truth at different noise levels of the observations. Our EnKG approach dramatically outperforms all other methods. Qualitatively, we see in Figure 2 that EnKG give solutions which qualitatively preserve important features of the flow, while some methods completely fail (i.e., overly noisy outputs).
On the computational aspect, the Navier-Stokes forward model (which requires a PDE solve) is extremely expensive, as shown in Figure 3(a). As such, the number of calls to the forward model dominates the computational cost. We see in Table 2 that our EnKG approach actually makes the fewest calls to the forward model (since it uses statistical linearization rather than trying to numerically approximate the gradient or value function), and thus EnKG is also the most computationally efficient approach, as seen in Figure 3(b). The traditional Ensemble Kalman Inversion (EKI) approach also employs statistical linearization, and so we do a detailed comparison in Figure 3(c), where we see that EnKG dominates EKI in the computational cost versus error trade-off curve.
Ablation study
For practical insights into selecting hyperparameters for EnKG, we perform ablation studies on the ensemble size, guidance scale, and the number of diffusion model queries. The results are shown in Figure 6 and Figure 8. We observe a clear trend: increasing the ensemble size consistently leads to improved performance, as measured by the relative $L^{2}$ error. However, the gains become marginal beyond 2048 particles. Our experiments also reveal that a guidance scale of 2.0 yields the best average performance across our experiments. Increasing the guidance scale beyond this point introduces instability, resulting in higher relative $L^{2}$ errors. This instability is likely due to the fact that large guidance scale might lead to divergence as Lemma 1 only guarantees the convergence with small guidance scale. Furthermore, we vary the number of diffusion model queries using different number of ODE steps in the likelihood estimation ($\hat{{\bm{x}}}_{N}^{(j)}$). We also examine the effect of varying the number of diffusion model queries by changing the number of ODE steps used in estimating $\hat{{\bm{x}}}_{N}^{(j)}$. Performance improves with more diffusion model queries but exhibits diminishing returns beyond 500 steps.
Figure 3: (a): runtime of single evaluation of the forward model, diffusion model, and diffusion model gradient (tested on a single A100). (b): comparison of computational characteristics of different algorithms on Navier-Stokes problem. Metrics are defined in the "Evaluation metrics" paragraph of Sec. 4.2. Each axis is normalized by dividing by the maximum over the algorithms. (c): compare EnKG with EKI on compute cost versus error.